# Cuff-less Arterial Blood Pressure Waveform Synthesis from Single-site PPG using Transformer & Frequency-domain Learning

### v1

@@ V-A Performance evaluation of Transformer model
Setting of hyperparameters: We used Adam optimizer, and used a learning rate of 1e-3. Further, we set segment size to 2 seconds, i.e., 250 samples. For training and evaluation of transformer model, we utilized an 80-20 train-test split. In multi-head attention layer, we set number of heads to 8, with the size of each attention head for query and key being 64. Next, the feedforward network consisted of 64 neurons and a relu activation function. The number of transformer layers was varied between 7-15. For layer normalization, we set epsilon to 1e-6. Finally, we set the dropout ratio to 0.1.
Fig. 3 is a qualitative result that showcases two examples where the transformer model synthesizes the ABP signal from a PPG signal. We make the following observations: 1) It is clear that the morphology of the two synthetic ABP waveforms constructed by the transformer model (on the right) is quite similar to the morphology of the reference ABP waveforms (on the left). 2) We note that the synthetic ABP waveform is shifted in amplitude, compared to the reference ABP waveform. This is the root cause of the error when we measure the SBP and DBP later by sampling the maximum and minimum point of the synthetic ABP waveform.
Now, the quantitative results. The lowest MAE (for ABP waveform synthesis) achieved by the transformer model is 9.10, with 15 layers. That is, the transformer model’s performance increases with increase in number of layers, and reaches a maximum for 15 layers (as further increasing the depth does not result in significant improvements). Further, the transformer model achieved an MAE of 19 and 9, for DBP and SBP, respectively.
Figure 3: Two examples of ABP waveform synthesis by our transformer model.
@@ V-B Performance evaluation of FD L/NL Regression model
Setting of hyperparameters: For ridge regression, we set $\alpha=0.5$. The rest of the hyperparameters remain the same.
Fig. 4 is again a qualitative result that demonstrates by means of two examples the quality of the ABP signal synthesized by the FD L/NL regression model. It is again evident that the morphology of the two synthetic ABP waveforms produced by the FD L/NL regression model (on the right) and the morphology of the reference ABP waveforms (on the left) are highly similar. Also, the synthetic ABP waveform is again shifted in amplitude, compared to the reference ABP waveform. But luckily, this time this shift in amplitude of the synthetic ABP is less pronounced. This helps us get a better MAE performance when we measure the SBP and DBP.
For our FD L/NL regression model, after converting the PPG signal to frequency-domain by means of the DCT, we have implemented ridge regression for linear regression purpose, and a feedforward neural network (FFNN) and a recurrent neural network (RNN) for non-linear regression purposes. The resulting MAE performance (for ABP waveform synthesis) is summarized in Table I. We learn from Table I that the ridge regression outperforms the FFNN and RNN by a large margin, which indicates that the relationship between the PPG signal and the ABP signal is perhaps linear.
Finally, our FD L/NL regression model (with ridge regression) achieves an MAE of 8.01 mmHg for DBP and an MAE of 11.87 mmHg for SBP.
TABLE I: Performance of different architectures for our FD L/NL regression model
FD L/NL regression model MAE
DCT + Ridge Regression 8.1
DCT + RNN 17.23
DCT + FFNN 20.24
@@ V-C Performance comparison with related work
Table II provides a detailed performance comparison of our work with the selected related works by providing a quick summary of different deep learning models used for CLBP estimation, their corresponding datasets, and the performance achieved by each model. Table II efficiently highlights the variability in the performance of the related works, which is in turn due to the change of data distribution across datasets, variable sizes of the datasets, model architectures, etc. Our first approach, i.e., the transformer based model achieved an MAE of 9.10 for ABP waveform synthesis which is at par with the related works. However, the transformer model achieved a relatively large MAE of 19, for DBP estimation, which suggests that our model is well-trained for ABP synthesis and SBP estimation only. Our 2nd approach, i.e., the FD L/NL regression model achieved a MAE of 8.01 for DBP and 11.87 for SBP, which is again at par with the performance reported by the related works.
TABLE II: Performance comparison of our work with the related work: SBP MAE and DBP MAE are in mmHg. $N$ represents number of subjects. * represents RMSE in mmHg.
Work
Dataset
$N$
Algorithm
SBP
DBP
Tazarv et al. [31]
MIMIC II
20
CNN-LSTM
3.70
2.02
Paviglianiti et al. [33]
MIMIC
40
ResNet, LSTM, WavNet, ResNet+LS-TM
4.118
2.228
Treebup-achatsa-kul et al.
UCI
812
CAN
7.14*
6.08*
Slapnicar et al.
MIMIC III
510
Spectrotem-poral ResNet
9.43
6.88
Wang et al.
UCI
348
Pretrained AlexNet, Inception-V3, VGG-19
6.17
3.16
Aguirre et al.
MIMIC
1131
RNN Encoder-Decoder + attention
14.39
14.39
Our work (transformer)
UCI
209
Transformer
19.0
9.0
Our work (FD L/NL regression)
UCI
209
DCT+L/NL Regression
11.87
8.01
TABLE III: Performance comparison of our work with related work according to the BHS criterion
Work Grade Cumulative Error
$\leq 5\%$ $\leq 10\%$ $\leq 15\%$
BHS A 60% 85% 95%
B 50 75 90
C 40 65 85
N. Ibtehaz et. al [18] A 87.4 95.2 97.7%
Tazarv et. al [31] A 75 92 96%
Li Y-H et. al B 59.46 79.97 88.45%
Kauchee et. al [8] C 44.7 71.6 86.7%
Yang et. al C 42.1 94.7 100%
Our work (FD L/NL regression) B 51.69 76.79 100%
Our work (transformer) C 36.47 71.98 100%
Finally, Table III outlines the performance grades of our transformer model and FD L/NL regression model, according to the BHS grading criterion. Specifically, Table III shows the cumulative percentages (CP) of the error achieved by the two models on the UCI test data. We can see that our transformer model achieves the grade C, while our FD L/NL regression model achieves the grade B. For the sake of completeness, Table III also summarizes the BHS performance grades obtained by the selected related works, for comparison purposes. We also want to highlight that nearly all the works suffer from the issue of lack of generalization, i.e., as the number of subjects increases, the MAE of the SBP and DBP increases, while the BHS grade tends to fall. On a side note, it is also worth mentioning that both of our models models fulfill the AAMI criterion of utilizing more than 85 subjects for training purposes.
@@ VI Conclusion
We evaluated two novel purpose-built DL models on public UCI dataset for CLBP estimation using a single-site PPG signal. The transformer model (the FD L/NL regression model) achieved an MAE of 19 and 9 (11.87 and 8.01), for DBP and SBP, respectively. Our FD L/NL regression model also fulfilled the AAMI criterion of utilizing data from more than 85 subjects, and achieved grade B by the BHS criterion.
Now, some parting thoughts. The UCI dataset consists of data from patients in intensive care units, who are often under the effects of drugs and other medical interventions. Thus, it is difficult to generalize the results of studies–that use the UCI dataset–to other populations, such as healthy adults. Thus, there is a need to evaluate our DL models on larger and diverse datasets in order to increase their generalization capability.

### v2

@@ V-A Performance evaluation of Transformer model
Setting of hyperparameters: We have a total of 32,052 samples/examples where each sample has duration 4 seconds (or, 1,000 data points). We set segment size to 2 seconds, i.e., 250 points. This changes the data shape from (32,052, 1,000) to (128,208, 250). For training and evaluation of transformer model, we utilized a 72:8:20 train-validation-test split. We set num-heads = 4, ff-dim = 64, num-transformer-blocks = 3, embed-dim = 64, epochs=3, batch-size=128, and optimizer=Adam. For each transformer block, we set epsilon to 1e-6 for normalization layer, and dropout ratio to 0.1.
Qualitative results: Fig. 2 shows four examples where the transformer model synthesizes the ABP signal from a PPG signal. It is clear that the morphology of the synthetic ABP waveforms constructed by the transformer model is quite similar to the morphology of the reference ABP waveforms.
Quantitative results: The transformer model achieves an MAE of 3.01 for ABP waveform synthesis, an MAE of 3.77 mmHg for SBP estimation, and an MAE of 2.69 mmHg for DBP estimation.
Figure 2: Four examples of ABP waveform synthesis by our transformer model.
Computational Complexity of Transformer Model: Recall that our transformer model consists of dense layers, positional encoding, transformer blocks, and multi-head attention block. We calculate its computational complexity by analyzing each layer and summing up the contributions.
[…]
  • Positional encoding: Positional encoding does not have parameters, and is computationally negligible.
  • Transformer blocks: A transformer block consists of multi-head self-attention and feed-forward network. Self-attention complexity: $O(N\times d^{2})$. Feed-forward network complexity: $O(N\times d^{2})$. In our work, $N=250,\ d=64$. This implies that the complexity of self-attention is: $O(250\times 64^{2})=O(1,024,000)$. The complexity of feed-forward network is: $O(250\times 64^{2})=O(1,024,000)$. Thus, total complexity for one transformer block is: $O(2\times 1,024,000)=O(2,048,000)$. Finally, total complexity for six transformer blocks is: $6\times 2,048,000=12,288,000$.
[…]
  • Overall complexity: We sum up the complexities of four dense layers, six transformer blocks, and multi-head attention block to get the total computational complexity of our transformer model: $O(1,072,000+12,288,000+8,192,000)=O(21,552,000)$.
@@ V-B Performance evaluation of FD L/NL Regression model
Setting of hyperparameters: For ridge regression, we vary the regularization parameter $\alpha$ in the range $31e-$ 3 to $31e$ 3. The optimal $\alpha$ turned out to be $01e$ 0. The rest of the hyperparameters remain the same.
Qualitative results: Fig. 3 presents two examples to showcase the quality of the ABP signal synthesized by the FD L/NL regression model. It is evident that the morphology of the two synthetic ABP waveforms produced by the FD L/NL regression model (on the right) and the morphology of the reference ABP waveforms (on the left) are highly similar.
Quantitative results: The FD L/NL model achieves an MAE of 4.23 for ABP waveform synthesis, an MAE of 4.37 mmHg for SBP estimation, and an MAE of 3.91 mmHg for DBP estimation.
@@
Computational Complexity of the FD L/NL Model: The computational complexity of the proposed FD L/NL model that involves taking the DCT of a PPG signal, computing the ridge regression of this DCT with the DCT of ABP signal, and then taking the I-DCT is given by:
  • DCT complexity: The complexity of computing the DCT of a signal of length $N$ is: $O(N\log N)$.
  • Ridge regression complexity: Assuming we have $m$ features and $n$ samples, the complexity of ridge regression is: $O(m^{2}n+m^{3})$.
  • I-DCT complexity: The complexity of computing the inverse DCT of a signal of length $N$ is: $O(N\log N)$.
  • Combined complexity: Combining the complexities of DCT, ridge regression, and IDCT, we get: $2\times O(N\log N)+O(m^{2}N+m^{3})$. In Big-O notation, constants are omitted, so the final computational complexity is: $O(N\log N)+O(m^{2}N+m^{3})$.
@@ V-C Performance comparison with related work
Table II provides a detailed performance comparison of our work with the selected related works by providing a quick summary of different deep learning models used for CLBP estimation, their corresponding datasets, MAE performance achieved by each model, and grade obtained according to the BHS and AAMI criteria. Table II highlights the variability in the performance of the related works, which is in turn due to the change of data distribution across datasets, variable sizes of the datasets, model architectures, etc. Our first approach, i.e., the transformer model achieves an MAE of 3.01 for ABP waveform synthesis, an MAE of 3.77 mmHg for SBP estimation, and an MAE of 2.69 mmHg for DBP estimation. Further, it fulfills AAMI criterion and obtains grade A by BHS criterion. Our 2nd approach, i.e., the FD L/NL regression model achieves an MAE of 4.23 for ABP waveform synthesis, an MAE of 4.37 mmHg for SBP estimation, and an MAE of 3.91 mmHg for DBP estimation. It also fulfills AAMI criterion and obtains grade B by BHS criterion. We note that nearly all the works suffer from the issue of lack of generalization, i.e., as the number of subjects increases, the MAE of the SBP and DBP increases, while the BHS grade tends to fall [40], [41].
TABLE II: Performance comparison of our work with the state-of-the-art CLBP estimation methods
Methodology MAE (SBP) MAE (DBP) Dataset Subjects BHS Grade AAMI Grade Reference
XGBoost 5.45 mmHg 3.88 mmHg custom SPG dataset 35 - - [38]
Waveform-based ANN-LSTM 2.34 mmHg 1.77 mmHg PPG-DaLiA 15 A Pass [42]
CNN with transfer learning 5.63 mmHg 2.82 mmHg MIMIC-II Database 942 A Pass [40]
DCT and Regression-Based Methods 4.02 mmHg 2.78 mmHg MIMIC Database 200 - Pass [41]
Stacked Deformable CNN 4.86 mmHg 4.42 mmHg Custom Clinical Dataset 30 A - [43]
CNN-LSTM 3.7 mmHg 2.02 mmHg MIMIC-II Dataset 20 A - [31]
This work (transformer method) 3.77 mmHg 2.69 mmHg UCI Database 209 A Pass -
This work (FD L/NL method) 4.37 mmHg 3.91 mmHg UCI Database 209 B Pass -
@@ VI Conclusion
We evaluated two novel purpose-built DL models on public UCI dataset for CLBP estimation using a single-site PPG signal. The transformer model (the FD L/NL regression model) achieved an MAE of 3.77 and 2.69 (4.37 and 3.91), for DBP and SBP, respectively. Both models fulfilled the AAMI criterion. Finally, the transformer model (FD L/NL model) achieved grade A (grade B), according to the BHS criterion. Thus, this work contributes to the efforts which aim to do SBP and DBP estimation using a single bio-signal obtained from a single site, e.g., through a smartwatch, smartband, etc.
We note that the UCI dataset consists of data from patients in intensive care units, who are often under the effects of drugs and other medical interventions. Thus, it is imperative to look into other populations, such as healthy adults doing various physical activities. Therefore, the future work will look into the subject-specific training and evaluation of our AI models, and evaluation and fine-tuning of our AI models on other diverse datasets to assess and enhance their generalization capability.